OpenStudy (anonymous):

how do i integrate (1+x^2)/(1-x^2) from 0 to 0.5

OpenStudy (zarkon):

expand (1+x^2)/(1-x^2)

OpenStudy (turingtest):

I wrote\[\int{1\over1-x^2}dx-\int{x^2\over1-x^2}dx\]the first one is a known hyperbolic tangent integral isn't it? I didn't want to say because I wasn't sure

OpenStudy (turingtest):

+ in between*

OpenStudy (turingtest):

the second is a trig sub for sure though...

OpenStudy (turingtest):

or long division...

OpenStudy (zarkon):

\[\frac{1+x^2}{1-x^2}=\frac{1}{x+1}-\frac{1}{x-1}-1\]

OpenStudy (turingtest):

oh yeah, that thing they call partial fractions !! I guess it eluded me at the moment :(

OpenStudy (anonymous):

another way i saw of doing it would be to write 2-(1-x^2)/ (1-x^2)

OpenStudy (anonymous):

[2-(1-x^2)]/ (1-x^2)

OpenStudy (turingtest):

still leads to partial fractions that way ... actually that is what I meant by long division

OpenStudy (turingtest):

...sort of

OpenStudy (anonymous):

yeah i guess. thanks for the help

OpenStudy (amistre64):

1+2x^2+2x^4+2x^6 ..... ------------- 1-x^2 ) 1+x^2 (1-x^2) -------- 2x^2 (2x^2-2x^4) ------------ 2x^4 .......... \[\int 1+2x^2+2x^4+2x^6 ...\ dx\] \[x+\frac{2}{3}x^3+\frac{2}{5}x^5+\frac{2}{7}x^7 ...\]

OpenStudy (amistre64):

\[\sum_{n=0}^{inf}\frac{2(0.5)^{2n+1}}{2n+1}\]